# Using Active Learning to Improve Quasar Identification for the DESI Spectra Processing Pipeline

### v1

@@ 4.2.1 Classification and Consistency Results
In Figure 7 we plot the estimated confusion values for each weights file. For our purposes we use the entropy as defined in Eq. 3.2 to determine confusion, as it is symmetric, i.e. the entropy value when 60% of the exposures classified as a quasar and 40% are not is the same as 40% being quasars and 60% being not. Since in this case there are only two possible classifications the maximum confusion value is $-\log_{2}{0.5}=1$ when the classifications are equally split between being a quasar and not. A confusion of 0 is only achieved when QuasarNET produces the same classification on all exposures. The left panel is a two dimensional plot of confusion values, with the vertical axis being the confusion from the DESI weights and the horizontal axis being the confusion from the eBOSS weights, demonstrating where confusion values lie in this two dimensional space. The right panel overplots 1-D histograms of the two weights-files’ results, underscoring the decrease in confusion values across all non-zero values with the new DESI weights. There is a corresponding slight increase in 0 values in the new DESI weights, but the overwhelming number of repeat exposures in that bin obfuscate this somewhat (note the logarithmic scaling on the vertical axis).
Figure 8: The mean binned standard deviation in redshift estimate. For each object we compute the standard deviation in redshift estimates from multiple exposures of the same object. Then we bin the objects in redshift, plotting the mean redshift of each bin and the mean of the standard deviations of the objects in that bin. In blue are the results from the prior eBOSS weights, and in orange are the results of the DESI weights. All bins above $z=4$ have low numbers of objects and corresponding higher variability. In all bins except one we produce more consistent redshift estimates.
In Figure 8 we show the results of the redshift consistency test, plotting the standard deviation of redshift estimates of the same object versus the redshift. For clarity we create bins in redshift, represented in the plot by their mean redshift value, and take the mean of the standard deviation values in that bin. In blue are the eBOSS weight’s redshift estimates, while below and in orange are the DESI weight’s estimates. The DESI weights produce consistently lower variance in redshift estimates, representing an improvement in consistency of redshift estimates.
@@ 4.2.2 Redshift Results and Identifying an Issue with QuasarNET
Some off-diagonal redshift changes can be explained by “line-confusion,” which is when two different weights files estimate redshift based on the position of the same emission line, but predict that the observer-frame emission line corresponds to two different emission lines in the rest-frame. This swap in line classification leads to an easily predictable change in redshift estimation. Some clearly identifiable line confusions are plotted along diagonals in light, dashed, gray along with associated labels. The first line name in each label is the predicted associated rest-frame line from the eBOSS weights, where the second is the prediction from the DESI weights.
@@ 5 Conclusions
In order to accommodate the reduced amount of training data we modified the structure of QuasarNET by including dropout layers and two additional convolution blocks. Both of these changes improve the generalizabiltiy of QuasarNET when trained on a reduced training set, an inherent problem when using DESI data only for training. Since visual inspection of DESI data has only been carried out on small subsamples of the full dataset, we only have approximately 4600 spectra available to use for training, a significant decrease compared to the eBOSS weights file.
@@
When used on DESI data the new weights file has marginally better completeness with a significant improvement in purity, reducing the number of false positive quasar identifications. When comparing different exposures of the same object our new weights file produces more consistent classifications, more often classifying objects the same way in different exposures of the same object. For objects that are consistently classified as a quasar our new weights file also has a significantly lower spread in redshift estimates.

### v2

@@ 4.3.1 Classification and Consistency Results
We will use the entropy as defined in Eq. 3.2 with two classes (QSO and NOT QSO) to determine classification confusion. The confusion is symmetric, for example when 60% of the exposures classified as a quasar and 40% are not the confusion value is the same as when 40% are labeled quasars and 60% are not. The maximum confusion value is $\log_{2}{0.5}=1$ when the classifications are equally split between being a quasar and not. A confusion of 0 is only achieved when QuasarNET produces the same classification on all exposures. In Figure 8 we plot the estimated confusion values for each weights file. The left panel is a two dimensional plot of confusion values, with the vertical axis being the confusion from the new DESI weights (trained on DESI AL 2) and the horizontal axis being the confusion from the eBOSS weights (used previously in DESI DR1, trained on eBOSS DR12), demonstrating where confusion values lie in this two dimensional space. The right panel overplots 1-D histograms of the two weights-files’ results, underscoring the decrease in confusion values across all non-zero values with the new DESI weights. There is a corresponding slight increase in 0 values in the new DESI weights, but the overwhelming number of repeat exposures in that bin obfuscate this somewhat (note the logarithmic scaling on the vertical axis).
Figure 9: The mean binned standard deviation in redshift estimate. For each object we compute the standard deviation in redshift estimates from multiple exposures of the same object. Then we bin the objects in redshift, plotting the mean redshift of each bin and the mean of the standard deviations of the objects in that bin. In orange are the results from the prior eBOSS weights, and in red are the results of the DESI weights. All bins above $z=4$ have low numbers of objects and corresponding higher variability. In all bins except one we produce more consistent redshift estimates.
In Figure 9 we show the results of the redshift consistency test, plotting the standard deviation of redshift estimates of the same object versus the redshift. For clarity we create bins in redshift, represented in the plot by their mean redshift value, and take the mean of the standard deviation values in that bin. In orange are the eBOSS weight’s redshift estimates, while below and in red are the DESI weight’s estimates. The DESI weights produce consistently lower variance in redshift estimates, representing an improvement in consistency of redshift estimates.
@@ 4.3.2 Redshift Results and Identifying an Issue with QuasarNET
Table 1: Table showing the number of spectra that have a velocity difference of $\leq 10000$ km/s away from the labeled confusions on Figure 10.
eBOSS Line Prediction DESI Line Prediction Number of Spectra
MgII(2796) CIV(1548) 76
MgII(2796) LYA 5
CIV(1548) CIII(1909) 70
LYA CIV(1548) 89
LYA CIII(1909) 17
LYA MgII(2796) 42
MgII(2796) CIII(1909) 8
@@
Some off-diagonal redshift changes can be explained by “line-confusion,” which is when two different weights files estimate redshift based on the position of the same emission line, but predict that the observer-frame emission line corresponds to two different emission lines in the rest-frame. This swap in line classification leads to an easily predictable change in redshift estimation. Some clearly identifiable line confusions are plotted along diagonals in light, dashed, gray along with associated labels. The first line name in each label is the predicted associated rest-frame line from the eBOSS weights, where the second is the prediction from the DESI weights. A summary of how many spectra whose velocity difference is $\leq 10000$ away from the labeled confusion is shown in Table 1. Many of these line confusions are a result of QuasarNET using only a single emission line for redshift identification, and might be resolved if the spectrum shows a second emission line to resolve the confusion. In some cases QuasarNET recognizes a second emission line but has no way to use this information in its internal line identification. Note that the number of spectra that cluster around these confusions is comparatively small ($O(100)$) compared to the total number of spectra used to generate the plot ($O(3000000)$).
@@ 5 Conclusions
In order to accommodate the reduced amount of training data we modified the structure of QuasarNET by including dropout layers and two additional convolution blocks. Both of these changes improve the generalizability of QuasarNET when trained on a reduced training set, an inherent problem when using DESI data only for training. Since visual inspection of DESI data has only been carried out on small subsamples of the full dataset, we only have approximately 4600 spectra available to use for training, a significant decrease compared to the eBOSS weights file.
@@
We demonstrate that the active learning methodology shows notable self improvement, improving the networks trained on the successive datasets produced by the algorithm. We also demonstrate that the active learning algorithm is selecting new spectra for training that expand the space of objects covered by the training dataset. We use the final product weights file, trained after two active learning iterations, as a candidate new DESI weights file.
When used on DESI data with DESI classification parameters the new weights file has somewhat better completeness with a similar purity, reducing the number of false negative quasar identifications. When comparing different exposures of the same object our new weights file produces more consistent classifications, more often classifying objects the same way in different exposures of the same object. For objects that are consistently classified as a quasar our new weights file also has a significantly lower spread in redshift estimates.